### v1

# On Computational Limits and Provably Efficient Criteria of Visual Autoregressive Models: A Fine-Grained Complexity Analysis

On Computational Limits and Provably Efficient Criteria of Visual Autoregressive Models: A Fine-Grained Complexity Analysis
Yekun Ke keyekun0628@gmail.com. Independent Researcher.    Xiaoyu Li 7.xiaoyu.li@gmail.con. Independent Researcher.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhizhou Sha shazz20@mails.tsinghua.edu.cn. Tsinghua University.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
Recently, Visual Autoregressive ($\mathsf{VAR}$) Models introduced a groundbreaking advancement in the field of image generation, offering a scalable approach through a coarse-to-fine “next-scale prediction” paradigm. However, the state-of-the-art algorithm of $\mathsf{VAR}$ models in [Tian, Jiang, Yuan, Peng and Wang, NeurIPS 2024] takes $O(n^{4})$ time, which is computationally inefficient. In this work, we analyze the computational limits and efficiency criteria of $\mathsf{VAR}$ Models through a fine-grained complexity lens. Our key contribution is identifying the conditions under which $\mathsf{VAR}$ computations can achieve sub-quadratic time complexity.
Specifically, we establish a critical threshold for the norm of input matrices used in $\mathsf{VAR}$ attention mechanisms. Above this threshold, assuming the Strong Exponential Time Hypothesis ($\mathsf{SETH}$) from fine-grained complexity theory, a sub-quartic time algorithm for $\mathsf{VAR}$ models is impossible. To substantiate our theoretical findings, we present efficient constructions leveraging low-rank approximations that align with the derived criteria.
Formally, suppose that $n$ is the height and width of the last VQ code map in $\mathsf{VAR}$ models, $d$ is the hidden dimension, $R$ is the bound of the entries of the input matrices for attention calculations in $\mathsf{VAR}$ models. We present two results:
  • On the positive side, we show that when $d=O(\log n)$ and $R=o(\sqrt{\log n})$, there is an $O(n^{2+o(1)})$-time algorithm that approximates the output of $\mathsf{VAR}$ model up to $1/\operatorname{poly}(n)$ additive error.
  • On the negative side, we show that when $d=O(\log n)$ and $R=\Theta(\sqrt{\log n})$, assuming $\mathsf{SETH}$, it is impossible to approximate the output of $\mathsf{VAR}$ model up to $1/\operatorname{poly}(n)$ additive error in truly sub-quartic time $O(n^{4-\Omega(1)})$.
This work initiates the study of the computational efficiency of the $\mathsf{VAR}$ model from a theoretical perspective. Our technique will shed light on advancing scalable and efficient image generation in $\mathsf{VAR}$ frameworks.
1 Introduction
Visual generation technologies now underpin a broad array of applications, ranging from image enhancement [27, 19] and augmented reality [9] to medical diagnostics [3, 40, 29] and creative pursuits like game development [45, 11]. By translating text descriptions or other input into detailed and diverse visuals, these models are reshaping both how machines interpret images and how new visual content is created. Leading methods in the field include Variational AutoEncoders (VAE) [15], Generative Adversarial Networks (GAN) [20], and Diffusion models [21]. Their advancements in producing high-resolution, high-fidelity, and varied imagery have significantly broadened the scope of visual generation, driving improvements in realism, diversity, and overall quality.
The emergence of the Visual AutoRegressive model ($\operatorname{\mathsf{VAR}}$) [51] marks a notable paradigm shift in image generation. Rather than relying on conventional “next-token prediction”, the $\operatorname{\mathsf{VAR}}$ model introduces a coarse-to-fine “next-scale prediction” approach, enabling autoregressive transformers to more efficiently learn visual distributions and outperform diffusion-based alternatives. Moreover, the $\operatorname{\mathsf{VAR}}$ model demonstrates robust zero-shot capabilities in tasks like image inpainting and editing, underscoring its potential for advancing autoregressive models in visual generation.
Despite its demonstrated strengths, there remains a critical need to investigate the $\operatorname{\mathsf{VAR}}$ model’s computational limits and to design efficient algorithms. In [51], the authors report that the $\operatorname{\mathsf{VAR}}$ model has a computational cost of $O(n^{4})$, improving upon the $O(n^{6})$ complexity associated with earlier autoregressive (AR) methods, where $n$ is the height and width of the last (largest) VQ code map. In this work, we aim to investigate the computational limits and potential efficient algorithms of $\operatorname{\mathsf{VAR}}$ models. Specifically, we ask the following questions:
Can we perform the computations of $\operatorname{\mathsf{VAR}}$ models faster than $O(n^{4})$ time?
We answer this question affirmatively and summarize our contributions as follows.
  • Computational Limits: We analyze the computation of the $\operatorname{\mathsf{VAR}}$ models under the Strong Exponential Time Hypothesis. Let $R$ represent the upper bound of the elements in the input matrices used for attention calculations in $\operatorname{\mathsf{VAR}}$ models. We establish an upper bound criterion $R^{*}=\Theta(\sqrt{\log n})$. Crucially, only when $R$ is below this threshold, one can compute $\operatorname{\mathsf{VAR}}$ models in $O(n^{4-\Omega(1)})$ time (truly sub-quartic time).
  • Provably Efficient Criteria: We further show that when $R=o(\sqrt{\log n})$, it becomes possible to design an algorithm that approximates the $\operatorname{\mathsf{VAR}}$ model in almost quadratic time, specifically $O(n^{2+o(1)})$.
1.1 Our Results
Our first result shows that when $R\geq\Omega(\sqrt{\log n})$, it is impossible to design a truly sub-quartic time algorithm. Our results for the lower bound make use of the Strong Exponential Time Hypothesis ($\mathsf{SETH}$[26] from the area of fine-grained complexity regarding the time required to solve $k$-SAT.
Theorem 1.1 (Computational Limits of $\operatorname{\mathsf{VAR}}$ Models, informal version of Theorem 4.4).
Suppose $d=O(\log n)$ and $R=\Theta(\sqrt{\log n})$. Assuming $\mathsf{SETH}$, there is no algorithm that approximates the $\operatorname{\mathsf{VAR}}$ model up to $1/\operatorname{poly}(n)$ additive error in $O(n^{4-\Omega(1)})$ time.
Our second result shows that when $R$ is $o(\sqrt{\log n})$, an almost quadratic time algorithm exists:
Theorem 1.2 (Existence of Almost Quadratic Time Algorithm, informal version of Theorem 5.8).
Suppose $d=O(\log n)$ and $R=o(\sqrt{\log n})$. There is an algorithm that approximates the $\operatorname{\mathsf{VAR}}$ model up to $1/\operatorname{poly}(n)$ additive error in $O(n^{2+o(1)})$ time.
Roadmap.
Section 2 offers a summary of related work. In Section 3, we outline the mathematical formulation of both the $\operatorname{\mathsf{VAR}}$ model and its fast version and divide the model into three stages: the $\operatorname{\mathsf{VAR}}$ Transformer, the feature map reconstruction block, and the VQVAE-Decoder. Section 4 delves into analyzing the computation limits of the $\operatorname{\mathsf{VAR}}$ model. In Section 5, we examine the running time and error propagation for each block in the fast $\operatorname{\mathsf{VAR}}$ model and establish the conditions under which the model can be accelerated with proven efficiency. In Section 6, we discuss the potential impacts and future directions. In Section 7, we conclude our contributions.
2 Related Work
2.1 Visual Generation Models
Recent years have witnessed remarkable advancements in visual generation models, driven by progress in several prominent architectures.
AutoRegressive Models.
AutoRegressive models for visual generation [16, 17] transform 2D images into 1D token sequences for processing. Early works like PixelCNN [52] and PixelSNAIL [14] pioneered pixel-by-pixel generation using a raster-scan approach. Subsequent studies [46, 18, 28] extended this concept by generating image tokens in a similar raster order. For example, VQ-GAN [18] employs a GPT-2-style decoder-only transformer for image generation, while models such as VQVAE-2 [46] and RQ-Transformer [28] enhance this method with additional hierarchical scales or stacked representations. More recently, Visual AutoRegressive ($\operatorname{\mathsf{VAR}}$) modeling [51] introduced a novel coarse-to-fine “next-scale prediction” approach. This method improves scalability, inference speed, and image quality, outperforming traditional autoregressive techniques and diffusion transformers.
Diffusion Models.
Diffusion models [21, 44] are known for their ability to generate high-resolution images by progressively refining noise into coherent visuals. Models such as DiT [43] and U-ViT [10] exemplify this approach, leveraging probabilistic frameworks to capture underlying data distributions. Recent advancements in diffusion-based generation focus on improving sampling techniques and training efficiency [47, 48, 38, 23], exploring latent-space learning [44, 56, 57, 39], enhancing model architectures [22, 43, 35, 53, 60], and 3D generation [42, 55, 59].
2.2 Acceleration via Low-rank Approximation
Low-rank approximation has emerged as a powerful technique for addressing the computational challenges associated with modern transformer architectures. By approximating key operations such as attention and gradient computations, these methods significantly reduce the time and resource requirements of training and inference.
Accelerating Attention Mechanisms.
Due to its quadratic computational complexity with respect to context length, the attention mechanism faces increasing difficulty as the sequence length grows in modern large language models [41, 2, 4]. To tackle this problem, polynomial kernel approximation methods [1] have been proposed, leveraging low-rank approximations to construct an efficient approximation of the attention matrix. These approaches lead to notable improvements in computation speed, enabling a single attention layer to handle both training and inference tasks with near-linear time complexity [5, 7]. Additionally, these methods can be extended to more sophisticated attention mechanisms, like tensor attention, while maintaining almost linear time complexity in both training and inference phases [8]. Furthermore, there are works considering RoPE-based attention [6, 12], and differentially private cross attention [36]. Additionally, alternative approaches like the conv-basis method introduced in [30] offer further opportunities for accelerating attention computations, providing complementary solutions to this critical bottleneck. Furthermore, there are many other works that use pruning to accelerate attention mechanisms [31, 13, 32, 50, 49, 25, 54, 58].
Gradient Approximation.
The low-rank approximation is a widely used technique for optimizing transformer training by reducing computational complexity [33, 37, 7, 24, 13, 34]. Specifically, [7] builds upon the low-rank approximation framework introduced in [5], which originally focused on forward attention computation, to approximate the gradient of attention mechanisms. This method effectively reduces the computational cost associated with gradient calculations. In [33], this low-rank gradient approximation approach is further extended to multi-layer transformers, demonstrating that backward computations in such architectures can be approximated in nearly linear time. Additionally, [37] generalizes the work of [7] to a tensor-based attention model, leveraging the forward computation results from [8] to enable efficient training of tensorized attention mechanisms. Finally, [24] utilizes low-rank approximation methods in the training process of Diffusion Transformers (DiTs)., highlighting the versatility of these methods in various transformer-based models.
3 Model Formulation
In Section 3.1, we give some definitions which will be used later. In Section 3.2, we present the mathematical formulation for the token map generation phase. In Section 3.3, we provide the mathematical formulation for the feature map reconstruction phase. Finally, in Section 3.4, we detail the mathematical formulation of the VQ-VAE Decoder within the $\operatorname{\mathsf{VAR}}$ model.
3.1 Notations and Definitions
We give the following notations in our setting. Given an integer $n\in\mathbb{Z}^{+}\cup\{0\}$, the set $\{1,2,\dots,n\}$ is represented by $[n]$. In our paper, nearly linear time is defined as $O(n\operatorname{poly}\log n)$, and almost linear time is defined as $O(n^{1+o(1)})$. Given a vector $c$, the diagonal matrix formed from $c$ is denoted as $\operatorname{diag}(c)$, where $c_{i}$ is the $i$-th diagonal entry of this matrix. Given a matrix $U$, we use $U^{\top}$ to denote the transpose of $U$. Given two vectors $a$ and $b$, which have the same length. The element-wise multiplication of $c$ and $d$ is denoted as $c\circ d$ with $i$-th entry being $c_{i}d_{i}$. Given a matrix $U$, we use $\|U\|_{F}$ to represent the Frobenius norm of $U$. Specifically, we have $\|U\|_{F}:=\sqrt{\sum_{i,j}U_{i,j}^{2}}$. Given a matrix $U$, we use $\|U\|_{\infty}$ to represent the maximum norm of $U$. Specifically, we have $\|U\|_{\infty}:=\max_{i,j}|U_{i,j}|$.
3.2 Phase 1: Token Maps Generation Phase
The $\operatorname{\mathsf{VAR}}$ model uses the $\operatorname{\mathsf{VAR}}$ Transformer to convert the initial tokens of the generated image into several pyramid-shaped token maps. And in the token maps generation phase, the token maps for the next scale, $M_{k+1}$, are generated based on the previous $k$ token maps ${M_{1},\dots,M_{k}}$. This phase has the main modules as the following:
Up Sample Blocks.
Before inputting the $k$ token maps into the $\operatorname{\mathsf{VAR}}$ Transformer, the $\operatorname{\mathsf{VAR}}$ model upsamples the $(i)$-th token map to the size of the $(i+1)$-th token map. Specifically, the $\operatorname{\mathsf{VAR}}$ model employs an upsampling method using interpolation for the image. Here, we define the up-interpolation layers:
Definition 3.1 (Bicubic Spline Kernel).
A bicubic spline kernel is a piecewise cubic function $W:\mathbb{R}\to\mathbb{R}$ that satisfies $W(x)\in[0,1]$ for all $x\in\mathbb{R}$.
Definition 3.2 (Up-interpolation Layer).
The Up-Interpolation layer is defined as follows:
  • Let $h\in\mathbb{N}$ and $h^{\prime}\in\mathbb{N}$ denote the height of the input feature map and output feature map, respectively.
  • Let $w\in\mathbb{N}$ and $w^{\prime}\in\mathbb{N}$ denote the width of the input feature map and the output feature map, respectively.
  • Let $c\in\mathbb{N}$ denote the number of channels of the input feature map and the output feature map.
  • Let $X\in\mathbb{R}^{h\times w\times c}$ denote the input feature map.
  • Let $Y\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$ denote the output feature map.
  • Let $s,t\in\{-1,0,1,2\}$.
  • Let $W:\mathbb{R}\to\mathbb{R}$ be a bicubic spline kernel as defined in 3.1.
    We use $\phi_{\rm up}:\mathbb{R}^{h\times w\times c}\to\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$ to denote the up-interpolation operation then we have $Y=\phi_{\rm up}(X)$. Specifically, for $i\in[h^{\prime}],j\in[w^{\prime}],l\in[c]$, we have
    $Y_{i,j,l}:=\sum_{s=-1}^{2}\sum_{t=-1}^{2}W(s)\cdot X_{\frac{ih}{h^{\prime}}+s,\frac{jw}{w^{\prime}}+t,l}\cdot W(t)$
Transformer Blocks.
After the up-sample process, the generated token maps above will be input into the Transformer to predict the next token map. Here, we define several blocks for the $\operatorname{\mathsf{VAR}}$ Transformer.
We first define the attention matrix.
Definition 3.3 (Attention Matrix).
Define the model weights as $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$, and let $X\in\mathbb{R}^{n\times d}$ represent the input of length $n$. The attention matrix $A\in\mathbb{R}^{n\times n}$ is given by the following, for $i,j\in[n],$
$A_{i,j}:=$ $~{}\exp(X_{i,*}W_{Q}W_{K}^{\top}X_{j,*}^{\top}.$
To move on, we define the single attention layer in the following way:
Definition 3.4 (Single Attention Layer).
Let $X\in\mathbb{R}^{n\times d}$ be the input matrix. Let $W_{V}\in\mathbb{R}^{d\times d}$ represent the weighted matrix of value. Similar to the standard attention mechanism, the goal is to produce an $n\times d$ output matrix, where $D:=\operatorname{diag}(A{\bf 1}_{n})\in\mathbb{R}^{n\times n}$. We define attention layer $\mathsf{Attn}$ as the following:
$\mathsf{Attn}(X):=$ $~{}D^{-1}AXW_{V}.$
Firstly, we give the definition of the $\operatorname{\mathsf{VAR}}$ Transformer Layer.
Definition 3.5 ($\operatorname{\mathsf{VAR}}$ Transformer Layer).
Given that the following conditions are true:
  • Define $X\in\mathbb{R}^{n\times d}$ as the input data matrix.
  • Let the up-interpolation layer $\phi_{\mathrm{up}}$ as given in Definition 3.2.
  • Let the $\mathsf{Attn}(X)$ as given in Definition 3.3.
Then, we defined one $\operatorname{\mathsf{VAR}}$ Transformer Layer as
$\mathsf{F}_{\mathrm{var}}(X):=\mathsf{Attn}(\phi_{\rm up}(X))$
Then, we can give the definition of the Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer.
Definition 3.6 (Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer).
Given that the following conditions are true:
  • Let $X\in\mathbb{R}^{n\times d}$ denote the input data matrix.
  • Let the up-interpolation layer $\phi_{\rm up}$ as given in Definition 3.2.
  • Let the $\mathsf{AAttC}(X)$ as given in Definition 4.2.
Then, we defined one Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer as
$\mathsf{F}_{\mathrm{fvar}}(X):=\mathsf{AAttC}(\phi_{\rm up}(X))$
Definition 3.7 ($\operatorname{\mathsf{VAR}}$ Model).
Given that the following conditions are true:
  • Let the initial input be defined as $x_{1}\in\mathbb{R}^{1\times d}$.
  • Let the $\operatorname{\mathsf{VAR}}$ Transformer layer $\mathsf{F}_{\mathrm{var}}$ as given in Definition 3.5.
  • Let $r\in\mathcal{N}$ represent the total iteration of $\operatorname{\mathsf{VAR}}$ model.
Then, for $i\in\{0,1,2,\cdots,r\}$, we defined the intermediate variable $T_{i}(X)$ as follows:
$T_{i}(X)=\begin{cases}x_{1},&~{}i=0\\ \mathsf{F}_{\mathrm{var}}(T_{i-1}(X)),&~{}i\in[r]\\ \end{cases}$
Then, the final output of the $\operatorname{\mathsf{VAR}}$ model is $T_{r}(X)$.
Definition 3.8 (Fast $\operatorname{\mathsf{VAR}}$ Model).
Given that the following conditions are true:
  • Define $x_{1}\in\mathbb{R}^{1\times d}$ as the initial input.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer layer $\mathsf{F}_{\mathrm{fvar}}$ be defined as Definition 3.6.
  • We take $R$ to be the upper bound on the entries of the input matrices involved in attention computations for $\operatorname{\mathsf{VAR}}$ models.
  • Let $r\in\mathcal{N}$ denote the total iteration of the Fast $\operatorname{\mathsf{VAR}}$ model.
Then, for $i\in\{0,1,2,\cdots,r\}$, we defined the intermediate variable $\tilde{T}_{i}(X)$ as follows:
$\widetilde{T}_{i}(X)=\begin{cases}x_{1},&~{}i=0\\ \mathsf{F}_{\mathrm{fvar}}(\widetilde{T}_{i-1}(X)),&~{}i\in[r]\\ \end{cases}$
Then, the final output of the $\operatorname{\mathsf{VAR}}$ model is $\tilde{T}_{r}(X)$.
3.3 Phase 2: Feature Map Reconstruction
In phase 2, the $\operatorname{\mathsf{VAR}}$ model will transform the generated token maps into feature maps. This phase has the following main modules:
Up Sample Blocks.
The $\operatorname{\mathsf{VAR}}$ model performs up-sampling on token maps of different sizes, scaling them to the size of the final output feature map. In this process, the $\operatorname{\mathsf{VAR}}$ model will use the up-interpolation blocks defined in Definition 3.2. To mitigate information loss during token map up-scaling, the $\operatorname{\mathsf{VAR}}$ model employs convolution blocks to post-process the up-scaled token maps. We define the convolution layers as the following:
Definition 3.9 (Convolution Layer).
The Convolution Layer is defined as follows:
  • Let $h\in\mathbb{N}$ denote the height of the input and output feature map.
  • Let $w\in\mathbb{N}$ denote the width of the input and output feature map.
  • Let $c_{\rm in}\in\mathbb{N}$ denote the number of channels of the input feature map.
  • Let $c_{\rm out}\in\mathbb{N}$ denote the number of channels of the output feature map.
  • Let $X\in\mathbb{R}^{h\times w\times c_{\rm in}}$ represent the input feature map.
  • For $l\in[c_{\rm out}]$, we use $K^{l}\in\mathbb{R}^{3\times 3\times c_{\rm in}}$ to denote the $l$-th convolution kernel.
  • Let $p=1$ denote the padding of the convolution layer.
  • Let $s=1$ denote the stride of the convolution kernel.
  • Let $Y\in\mathbb{R}^{h\times w\times c_{\rm out}}$ represent the output feature map.
We use $\phi_{\rm conv}:\mathbb{R}^{h\times w\times c_{\rm in}}\to\mathbb{R}^{h\times w\times c_{\rm out}}$ to represent the convolution operation then we have $Y=\phi_{\rm conv}(X)$. Specifically, for $i\in[h],j\in[w],l\in[c_{\rm out}]$, we have
$Y_{i,j,l}:=\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{c=1}^{c_{\rm in}}X_{i+m-1,j+n-1,c}\cdot K^{l}_{m,n,c}+b$
Remark 3.10.
Assumptions of kernel size, padding of the convolution layer, and stride of the convolution kernel are based on the specific implementation of [51].
3.4 Phase 3: VQ-VAE Decoder process
$\operatorname{\mathsf{VAR}}$ will use the VQ-VAE Decoder Module to reconstruct the feature map generated in Section 3.3 into a new image. The Decoder of VQ-VAE has the following main modules:
ResNet Blocks.
In the VQVAE decoder, the ResNet block, which includes two (or more) convolution blocks, plays a crucial role in improving the model’s ability to reconstruct high-quality outputs. The convolution blocks help capture spatial hierarchies and patterns in the data, while the residual connections facilitate better gradient flow and allow the model to focus on learning the residuals (differences) between the input and output. The definition of convolution block is given in Definition 3.9.
Attention Blocks.
The Attention block helps the Decoder fuse information from different locations during the generation process, which can significantly improve the clarity and detail of the generated images. When applied to a feature map, the attention mechanism computes attention scores for all pairs of pixels, capturing their pairwise relationships and dependencies. The definitions of blocks in attention are given in Section 3.2.
Up Sample Blocks.
The VQ-VAE decoder uses Up-Sample Blocks to progressively increase the spatial resolution of the latent representation. The Up-Sample Blocks in VQVAE combine up-interpolation and convolution blocks to restore the spatial dimensions of the feature maps, facilitating the reconstruction of the high-resolution output image. The convolution block has already been defined in Definition 3.9, and the up-interpolation block has already been defined in Definition 3.2.
4 Computational Limits
In this section, we delve into the computational limits of $\operatorname{\mathsf{VAR}}$ Models, particularly in the context of solving key problems under the assumptions of the Strong Exponential Time Hypothesis ($\mathsf{SETH}$). Section 4.1 introduces $\mathsf{SETH}$ as the basis for our complexity analysis. In Section 4.2, we discuss a key result from [5] that establishes the hardness of Approximate Attention Computation. Finally, Section 4.3 presents the lower bound for $\operatorname{\mathsf{VAR}}$ model efficiency, pinpointing the limitations for sub-quartic performance.
4.1 Strong Exponential Time Hypothesis
We begin by presenting the foundational hypothesis ($\mathsf{SETH}$[26], which underpins much of our complexity analysis:
Hypothesis 4.1 (Strong Exponential Time Hypothesis ($\mathsf{SETH}$[26]).
[…]
4.2 Hardness of Approximate Attention Computation
We begin by introducing the definition of Approximate Attention Computation ($\mathsf{AAttC}$).
Definition 4.2 (Approximate Attention Computation $\mathsf{AAttC}(n,d,B,\delta)$, Definition 1.2 in [5]).
Let $\delta>0$. Let $R>0$. Let $X\in\mathbb{R}^{n\times d}$ denote the input of the attention mechanism. Given three matrices $Q,K,V\in\mathbb{R}^{n\times d}$, with the guarantees that $\|Q\|_{\infty}\leq R$, $\|K\|_{\infty}\leq R$ and $\|V\|_{\infty}\leq R$, output a matrix $T\in\mathbb{R}^{n\times d}$ that approximately represents $\mathsf{Attn}(X)$, meaning
$\|T-\mathsf{Attn}(X)\|_{\infty}\leq\delta$
Next, we state a result for the Approximate Attention Computation (AAttC) from [5].
Lemma 4.3 (Theorem 4.6 in [5]).
Suppose $d=O(\log n)$ and $R=\Theta(\sqrt{\log n})$. Assuming $\mathsf{SETH}$, there is no algorithm that solves the Approximate Attention Computation $(\mathsf{AAttC})$ up to $1/\operatorname{poly}(n)$ additive error in $O(n^{4-\Omega(1)})$ time.
4.3 Computational Limits of Fast $\operatorname{\mathsf{VAR}}$ Models
We now present a main theorem detailing the lower bound for $\operatorname{\mathsf{VAR}}$ model computation.
Theorem 4.4 (Computational Limits of Fast $\operatorname{\mathsf{VAR}}$ Models, formal version of Theorem 1.1).
Suppose $d=O(\log n)$ and $R=\Theta(\sqrt{\log n})$. Assuming $\mathsf{SETH}$, there is no algorithm that approximates the $\operatorname{\mathsf{VAR}}$ model up to $1/\operatorname{poly}(n)$ additive error in $O(n^{4-\Omega(1)})$ time.
Proof.
By Lemma 4.3, in the $K$-th step ($K=\log_{a}n$), the $\operatorname{\mathsf{VAR}}$ Transformer must compute attention with a computational cost at least
$\Omega(L_{K}^{2-q}\cdot d)=$ $~{}\Omega((\sum_{i=1}^{K}(\alpha^{i-1})^{2})^{2-q}\cdot d)$
$=$ $~{}\Omega((\frac{\alpha^{2K}-1}{\alpha^{2}-1})^{2-q}\cdot d)$
$\geq$ $~{}\Omega((\alpha^{2K}-1)^{2-q}\cdot d)$
$\geq$ $~{}\Omega(n^{4-2q}\,d).$
In the first step above, we use the definition of $L_{K}$. The second step applies the standard geometric series formula. The third step involves basic algebra, and the final inequality is due to the fact $K=\log_{a}n$. ∎
5 Provably Efficient Criteria
Section 5.1 details the running time of the fast $\operatorname{\mathsf{VAR}}$ Transformer, feature map reconstruction block, and Fast VQ-VAE Decoder. In Section 5.2, we analyze the error propagation in both the Fast $\operatorname{\mathsf{VAR}}$ Transformer and the Fast VQ-VAE Decoder. Section 5.3 presents our findings regarding the existence of an almost quadratic time algorithm.
5.1 Running Time
Here, we present an overview of the computational cost associated with the Fast $\operatorname{\mathsf{VAR}}$ Transformer, feature map reconstruction block, and Fast VQ-VAE Decoder.
Firstly, we show that the runtime of the $\operatorname{\mathsf{VAR}}$ Transformer can be sped up to $O(n^{2+o(1)})$.
Lemma 5.1 (Running time of Fast $\operatorname{\mathsf{VAR}}$ Transformer, informal version of Lemma E.3).
Assuming the conditions below are satisfied:
  • Let $K$ denote the total number of the generated token maps.
  • Let $k\in[K]$.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k\in[K]$, the $k$-th token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$.
  • Let $r_{K}\in\mathbb{R}^{n\times n\times d}$ denote the last scale token map, where $n=\alpha^{K-1}$.
  • Let $d=O(\log(n))$.
Then, the total runtime of the $\operatorname{\mathsf{VAR}}$ Transformer for generating token maps can be accelerated to $O(n^{2+o(1)})$.
Then, we proceed to show that the runtime of the feature map reconstruction layer is $O(n^{2+o(1)})$.
Lemma 5.2 (Running time of Feature Map Reconstruction Layer, informal version of Lemma E.4).
Assuming the conditions below are satisfied:
  • Let $K$ denote the total number of generated token maps.
  • Let $k\in[K]$.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k\in[K]$, the $k$-th token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$.
  • Let $r_{K}\in\mathbb{R}^{n\times n\times d}$ denote the last scale token map, where $n=\alpha^{K-1}$.
  • Let $d=O(\log(n))$.
Then, the total runtime of the Feature Map Reconstruction Layer is $O(n^{2+o(1)})$.
Finally, we show that the runtime of the VQVAE Decoder can be sped up to $O(n^{2+o(1)})$.
Lemma 5.3 (Running time of Fast VQ-VAE Decoder, informal version of Lemma E.6).
Assuming the conditions below are satisfied:
  • Define $k_{1},k_{2},k_{3}\in\mathbb{N}$ as constant numbers.
  • Let $X\in\mathbb{R}^{n\times n\times d}$ represent the input feature map.
  • We assume there are $k_{1}$ up-interpolation layers $\phi_{\rm up}$ defined in Definition 3.2.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{O(h)\times O(w)\times d}$.
  • We assume there are $k_{2}$ attention layers $\mathsf{Attn}$ defined in Definition 3.4.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, the $i$-th attention layer’s output $\mathsf{Attn}(M)\in\mathbb{R}^{h\times w\times d}$.
  • We assume there are $k_{3}$ convolution layers $\phi_{\rm conv}$ defined in Definition 3.9.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
  • Let $d=O(\log(n))$.
Then, the total runtime of the VQ-VAE decoder can be accelerated to $O(n^{2+o(1)})$.
5.2 Error Propagation Analysis
Here, we present the error analysis results of the fast Transformer layer and the fast VQ-VAE Decoder.
We begin with analyzing the error propagation of $\mathsf{AAttC}$ (see Definition 4.2).
Lemma 5.4 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$, informal version of Lemma C.3).
Assuming the conditions below are satisfied:
  • Define $X\in\mathbb{R}^{n\times d}$ as the input matrix.
  • Define $X^{\prime}\in\mathbb{R}^{n\times d}$ as the approximation version of input matrix.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)$ as given in Definition 4.2.
  • Let $U,V\in\mathbb{R}^{n\times k}$ represent low-rank matrices constructed for polynomial approximation of attention matrix $\mathsf{AAttC}(X)$.
  • Let $f$ be a polynomial with degree $g$.
Then, we can show that
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X)\|_{\infty}\leq O(kM^{g+2}d)\cdot\epsilon$
Then, we move on to analyzing the approximation error between $\mathsf{AAttC}(X^{\prime})$ and the ground truth $\mathsf{Attn(X)}$, where $X^{\prime}$ denotes the approximated version of the original input $X$.
Lemma 5.5 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$, informal version of Lemma C.4).
Assuming the conditions below are satisfied:
  • Define $X\in\mathbb{R}^{n\times d}$ as the input matrix.
  • Define $X^{\prime}\in\mathbb{R}^{n\times d}$ as the approximation version of input matrix.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the attention matrix $\mathsf{Attn}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 3.3.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 4.2.
  • Let $U,V\in\mathbb{R}^{n\times k}$ be low-rank matrices constructed for polynomial approximation of attention matrix $\mathsf{AAttC}(X)$.
  • Let $f$ be a polynomial with degree $g$.
Then, we can show that
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{Attn}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
With the above foundations, we are ready to move to analyzing the approximation error of one fast $\operatorname{\mathsf{VAR}}$ layer (see Definition 3.6).
Lemma 5.6 (Error propagation analysis for one $\operatorname{\mathsf{VAR}}$ Transformer Layer, informal version of Lemma C.7).
Assuming the conditions below are satisfied:
  • Define $X\in\mathbb{R}^{n\times d}$ as the input data matrix.
  • Define $X^{\prime}\in\mathbb{R}^{n\times d}$ as the approximation version of $X$.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let the $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\mathrm{var}}$ as given in Definition 3.5.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\rm fvar}$ as given in Definition 3.6.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the attention matrix $\mathsf{Attn}(X)\in\mathbb{R}^{n\times n}$ be defined as Definition 3.3.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)\in\mathbb{R}^{n\times n}$ be defined as Definition 4.2.
  • Let $U,V\in\mathbb{R}^{n\times k}$ be low-rank matrices constructed for polynomial approximation of attention matrix $\mathsf{AAttC}(X)$.
  • Let $f$ be a polynomial with degree $g$.
Then, we can show that
$\|\mathsf{F}_{\rm fvar}(X^{\prime})-\mathsf{F}_{\mathrm{var}}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
Then, we show that the $\operatorname{\mathsf{VAR}}$ Transformer will introduce additive error bounded by $1/\operatorname{poly}(n)$.
Lemma 5.7 (Error analysis for the Fast $\operatorname{\mathsf{VAR}}$ Transformer, informal version of Lemma C.8).
Assuming the conditions below are satisfied:
  • Define $X_{0}\in\mathbb{R}^{1\times d}$ as the initial input data matrix.
  • Define $r\in\mathcal{N}$ as the total iteration of the Fast $\operatorname{\mathsf{VAR}}$ model.
  • For $i\in[r]$, let $T_{i}(X_{0})$ denote the intermediate variable of the $\operatorname{\mathsf{VAR}}$ model, as defined in Definition 3.7.
  • For $i\in[r]$, let $\widetilde{T}_{i}(X_{0})$ denote the intermediate variable of the fast $\operatorname{\mathsf{VAR}}$ model, as defined in Definition 3.8.
  • Assume $T_{r}(X_{0})\in\mathbb{R}^{O(n^{2})\times d}$ is final output of the $\operatorname{\mathsf{VAR}}$ Transformer.
  • Assume $\widetilde{T}_{r}(X_{0})\in\mathbb{R}^{O(n^{2})\times d}$ is final output of the Fast $\operatorname{\mathsf{VAR}}$ Transformer.
  • Assume each element of the matrices is encoded in $O(\log n)$ bits.
Then, we can show that the error bound of the final output $\widetilde{T}_{r}(X_{0})$ as
$\|\widetilde{T}_{r}(X_{0})-T_{r}(X_{0})\|_{\infty}\leq 1/\operatorname{poly}(n)$
5.3 Existence of Almost Quadratic Time Algorithm
This section presents a theorem proving the existence of a quadratic-time algorithm that speeds up the $\operatorname{\mathsf{VAR}}$ model and guarantees a bounded additive error.
Theorem 5.8 (Existence of Almost Quadratic Time Algorithm, formal version of Theorem 1.2).
Suppose $d=O(\log n)$ and $R=o(\sqrt{\log n})$. There is an algorithm that approximates the $\operatorname{\mathsf{VAR}}$ model up to $1/\operatorname{poly}(n)$ additive error in $O(n^{2+o(1)})$ time.
Proof.
By combining the result of Lemma 5.1, Lemma 5.2, Lemma 5.3 and Lemma 5.7, we can easily derive the proof. ∎
6 Discussion
The fine-grained analysis of Visual Autoregressive ($\operatorname{\mathsf{VAR}}$) models we provided in this paper uncovers the critical computational limitations and proposes criteria that ensure efficiency under the Strong Exponential Time Hypothesis ($\mathsf{SETH}$). The insights from this analysis are not only important for $\operatorname{\mathsf{VAR}}$ models but also carry broader implications for the deep learning and machine learning communities as a whole. One of the key contributions of this work is that understanding the computational bottlenecks of $\operatorname{\mathsf{VAR}}$ models allows us to more clearly delineate the theoretical boundaries of model performance, which in turn helps guide the design of future models.
By exploring the specific conditions under which the $\operatorname{\mathsf{VAR}}$ models hit computational limits, it is important to identify and address these bottlenecks early in the model development process. This understanding can prevent the misallocation of resources toward achieving computational feats that are not feasible, particularly in the context of autoregressive models used for visual generation tasks. In particular, demonstrating that sub-quartic time complexity is unattainable when input matrices exceed a critical threshold provides a crucial reference point for the deep learning community. This knowledge empowers researchers to set realistic expectations regarding model efficiency and to focus their efforts on optimizations that are computationally viable.
This work provides a foundational framework for understanding and overcoming the computational bottlenecks in generative models. It will serve as a key resource for researchers striving to design the next generation of efficient autoregressive models. By addressing the limitations of current models and offering clear guidance on how to optimize them, we hope to inspire more efficient and scalable solutions for a wide array of machine learning applications, extending far beyond visual generation.
7 Conclusion
This paper provides a fine-grained complexity analysis of Visual Autoregressive ($\operatorname{\mathsf{VAR}}$) models, identifying computational limits and efficient criteria under the Strong Exponential Time Hypothesis ($\mathsf{SETH}$). By rigorously analyzing computational trade-offs and proposing provably efficient criteria, this work establishes a foundational understanding that will guide the development of next-generation autoregressive models in visual generation. We demonstrate the infeasibility of achieving sub-quartic time complexity for $\operatorname{\mathsf{VAR}}$ computations when the norm of input matrices exceeds a critical threshold. In contrast, we establish that sub-quadratic time approximations become feasible under carefully designed conditions, leveraging low-rank approximations. In future works, we will explore the extension of these methods to other domains where autoregressive models play a pivotal role, such as text-to-image synthesis and multi-modal generation tasks. Additionally, integrating hardware acceleration strategies could further optimize the computational pipeline, broadening the applicability of $\operatorname{\mathsf{VAR}}$ models in resource-constrained environments.
Appendix
Roadmap. Section A introduces key notations. Section B provides an analysis of the perturbation error in different VAR model modules. Section C details the error analysis for the VAR Transformer. In Section D, we present the error analysis of the VQVAE decoder. In Section E, we evaluate the running time of VAR models and fast VAR models.
Appendix A Notations
Given an integer $n\in\mathbb{Z}^{+}\cup\{0\}$, the set $\{1,2,\dots,n\}$ is represented by $[n]$. In our paper, nearly linear time is defined as $O(n\operatorname{poly}\log n)$, and almost linear time is defined as $O(n^{1+o(1)})$. Given a vector $c$, the diagonal matrix formed from $c$ is denoted as $\operatorname{diag}(c)$, where $c_{i}$ is the $i$-th diagonal entry of this matrix. Given a matrix $U$, we use $U^{\top}$ to denote the transpose of $U$. Given two vectors $a$ and $b$, which have the same length. The element-wise multiplication of $c$ and $d$ is denoted as $c\circ d$ with $i$-th entry being $c_{i}d_{i}$. Given a matrix $U$, we use $\|U\|_{F}$ to represent the Frobenius norm of $U$. Specifically, we have $\|U\|_{F}:=\sqrt{\sum_{i,j}U_{i,j}^{2}}$. Given a matrix $U$, we use $\|U\|_{\infty}$ to represent the maximum norm of $U$. Specifically, we have $\|U\|_{\infty}:=\max_{i,j}|U_{i,j}|$.
Appendix B Perturbation Error Analysis
Section B.1 focuses on the perturbation error of linear interpolation. In Section B.2, we explore the error incurred when raising an approximated value to a power. Section B.3 analyzes the perturbation error of approximating the product of two values, while Section B.4 addresses the error in approximating the product of $d$ values.
B.1 Perturbation Error for Linear Interpolation
First, we state a fact that demonstrates that the simple linear interpolation operation does not introduce additional error.
Fact B.1.
Assuming the conditions below are satisfied:
  • Let $a,b\in\mathbb{R}$.
  • Define $a^{\prime},b^{\prime}\in\mathbb{R}$ as the approximated value of $a,b$.
  • Let $|a|\leq R$ and $|a^{\prime}|\leq R$.
  • Let $|b|\leq R$ and $|b^{\prime}|\leq R$.
  • Let $\delta\in(0,0.1)$ denote the approximation error.
  • Suppose $|a-a^{\prime}|\leq\delta$ and $|b-b^{\prime}|\leq\delta$.
Then we have, for any $w\in(0,1)$
$|(wa^{\prime}+(1-w)b^{\prime})-(wa+(1-w)b)|\leq\delta$
Proof.
We have
$~{}|(wa^{\prime}+(1-w)b^{\prime})-(wa+(1-w)b)|$
$=$ $~{}|w(a^{\prime}-a)+(1-w)(b^{\prime}-b)|$
$\leq$ $~{}w|a^{\prime}-a|+(1-w)|b^{\prime}-b|$
$\leq$ $~{}w\cdot\delta+(1-w)\cdot\delta$
$=$ $~{}\delta$
We derive the first step using basic algebra, the second step using the triangle inequality and the fact that $w\in(0,1)$, the third step using the conditions of the lemma, and the last step using basic algebra.
B.2 Perturbation Error of Power Function
We then present several lemmas that will be utilized in the subsequent proof.
Fact B.2.
Assuming the conditions below are satisfied:
  • Let $a\in\mathbb{R}$.
  • Let $a^{\prime}\in\mathbb{R}$ denote the approximation value of $a$.
  • Let $|a|\leq R$.
  • Let $|a^{\prime}|\leq R$.
  • Define $\delta\in(0,0.1)$ as the approximation error.
  • Suppose $\delta\leq|a|$.
  • Suppose $|a^{\prime}-a|\leq\delta$.
  • Let $m\in\mathbb{N}$.
  • Let $C:=(2^{m}-1)\cdot R^{m-1}$.
Then we have
$|(a^{\prime})^{m}-a^{m}|\leq C\cdot\delta$
Proof.
Case 1: We firstly consider the case $|(a^{\prime})^{m}-a^{m}|\leq|(a+\delta)^{m}-a^{m}|$. Then we have:
$|(a^{\prime})^{m}-a^{m}|\leq$ $~{}|(a+\delta)^{m}-a^{m}|$
$\leq$ $~{}|\sum_{i=0}^{m}\binom{m}{i}a^{m-i}\delta^{i}-a^{m}|$
$=$ $~{}|\sum_{i=1}^{m}\binom{m}{i}a^{m-i}\delta^{i}|$
$\leq$ $~{}\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}\delta^{i}|$
The first step above is based on the condition of this lemma, the second relies on binomial expansion, the third uses simple algebra, and the last step follows from the triangle inequality.
Then, we keep on showing that:
$\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}\delta^{i}|=$ $~{}\sum_{i=1}^{m}\binom{m}{i}\cdot|a^{m-i}|\cdot\delta^{i}$
$\leq$ $~{}\sum_{i=1}^{m}\binom{m}{i}\cdot|a^{m-1}|\cdot\delta$
$\leq$ $~{}(2^{m}-1)\cdot R^{m-1}\cdot\delta$
$=$ $~{}C\cdot\delta$
The first step above is based on $\delta>0$ and $\binom{m}{i}>0$. The second step is derived from the condition that $\delta\leq|a|$. The third step is due to the fact that $\sum_{i=1}^{m}\binom{m}{i}=2^{m}-1$ and the condition that $|a|\leq R$. The final step is trivially from the definition of $C$.
Case 2: Then we consider the case $|(a^{\prime})^{m}-a^{m}|\leq|(a-\delta)^{m}-a^{m}|$. Then we have:
$|(a^{\prime})^{m}-a^{m}|\leq$ $~{}|(a-\delta)^{m}-a^{m}|$
$\leq$ $~{}|\sum_{i=0}^{m}\binom{m}{i}a^{m-i}(-\delta)^{i}-a^{m}|$
$=$ $~{}|\sum_{i=1}^{m}\binom{m}{i}a^{m-i}(-\delta)^{i}|$
$\leq$ $~{}\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}(-\delta)^{i}|$
$=$ $~{}\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}\delta^{i}|$
The first step relies on this lemma’s condition, the second on binomial expansion, the third on basic algebra, the fourth on triangle inequality, and the last on basic algebra.
Then, we can further show that:
$\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}\delta^{i}|=$ $~{}\sum_{i=1}^{m}\binom{m}{i}\cdot|a^{m-i}|\cdot\delta^{i}$
$\leq$ $~{}\sum_{i=1}^{m}\binom{m}{i}\cdot|a^{m-1}|\cdot\delta$
$\leq$ $~{}(2^{m}-1)\cdot R^{m-1}\cdot\delta$
$=$ $~{}C\cdot\delta$
The first step above relies on the condition that $\delta>0$ and $\binom{m}{i}>0$. And the second step uses the condition that $\delta\leq|a|$. The third step derives from the fact that $\sum_{i=1}^{m}\binom{m}{i}=2^{m}-1$ and the condition that $|a|\leq R$, and the final step is from the definition of $C$.
B.3 Perturbation Error of Production of Two Terms
In this section, we introduce the perturbation error of the production of two terms.
Fact B.3.
Assuming the conditions below are satisfied:
  • Let $a,b\in\mathbb{R}$.
  • Define $a^{\prime},b^{\prime}\in\mathbb{R}$ as the approximated values of $a,b$.
  • Let $|a|\leq R$
  • Let $|a^{\prime}|\leq R$.
  • Let $|b|\leq R$
  • Let $|b^{\prime}|\leq R$.
  • Let $\delta\in(0,0.1)$ denote the approximation value.
  • Suppose we have $|a^{\prime}-a|\leq\delta$.
  • Suppose we have $|b^{\prime}-b|\leq\delta$.
Then we have
$|a^{\prime}b^{\prime}-ab|\leq 2R\cdot\delta$
Proof.
We have
$|a^{\prime}b^{\prime}-ab|=$ $~{}|(a^{\prime}b^{\prime}-a^{\prime}b)+(a^{\prime}b-ab)|$
$=$ $~{}|a^{\prime}(b^{\prime}-b)+b(a^{\prime}-a)|$
$\leq$ $~{}|a^{\prime}|\cdot\delta+|b|\cdot\delta$
$\leq$ $~{}2R\cdot\delta$
Starting with simple algebra for the first and second steps, the third step uses the triangle inequality along with $|a^{\prime}-a|\leq\delta$ and $|b-b^{\prime}|\leq\delta$, and the final step is based on $|a^{\prime}|\leq R$ and $b\leq R$. ∎
B.4 Perturbation Error of Production of $d$ Terms
In this section, we introduce the perturbation error of production of $d$ terms.
Fact B.4.
Assuming the conditions below are satisfied:
  • Let $a_{1},a_{2},\cdots,a_{d}\in\mathbb{R}$.
  • Define $a^{\prime}_{1},a^{\prime}_{2},\cdots,a^{\prime}_{d}\in\mathbb{R}$ as the approximation values of $a_{1},a_{2},\cdots,a_{d}$.
  • For every $i\in[d]$, we assume $|a_{i}|\leq R$ and $|a^{\prime}_{i}|\leq R$.
  • Let $\delta\in(0,0.1)$ denote the approximation error.
  • For every $i\in[d]$, we assume $|a_{i}-a^{\prime}_{i}|\leq\delta$.
Then we have
$|\prod_{i=1}^{d}a_{i}-\prod_{i=1}^{d}a^{\prime}_{i}|\leq d\cdot R^{d-1}\cdot\delta$
Proof.
The proof is carried out by mathematical induction as follows:
Consider the case $d=2$. We can show that
$|\prod_{i=1}^{2}a_{i}-\prod_{i=1}^{2}a^{\prime}_{i}|=$ $~{}|a_{1}a_{2}-a^{\prime}_{1}a^{\prime}_{2}|$
$\leq$ $~{}2R\cdot\delta$
The reasoning above follows from Fact B.3. Assume the statement is true for $d=k$:
$|\prod_{i=1}^{k}a_{i}-\prod_{i=1}^{k}a^{\prime}_{i}|\leq k\cdot R^{k-1}\cdot\delta$ (1)
consequently, we can demonstrate that:
$~{}|\prod_{i=1}^{k+1}a_{i}-\prod_{i=1}^{k+1}a^{\prime}_{i}|$
$=$ $~{}|(\prod_{i=1}^{k}a_{i})\cdot a_{k+1}-(\prod_{i=1}^{k}a^{\prime}_{i})\cdot a^{\prime}_{k+1}|$
$=$ $~{}|(\prod_{i=1}^{k}a_{i})\cdot(a_{k+1}-a^{\prime}_{k+1})+(\prod_{i=1}^{k}a_{i}-\prod_{i=1}^{k}a^{\prime}_{i})\cdot a^{\prime}_{k+1}|$
$\leq$ $~{}|(\prod_{i=1}^{k}a_{i})\cdot(a_{k+1}-a^{\prime}_{k+1})|+|(\prod_{i=1}^{k}a_{i}-\prod_{i=1}^{k}a^{\prime}_{i})\cdot a^{\prime}_{k+1}|$
$\leq$ $~{}R^{k}\cdot\delta+k\cdot R^{k-1}\cdot\delta\cdot R$
$=$ $~{}(k+1)\cdot R^{k}\cdot\delta$
Starting with simple algebra for the first and second steps, the third step follows from the triangle inequality, the fourth from the conditions of this lemma and Eq. (1), and the final step involves simple algebra.
Hence, we have
$|\prod_{i=1}^{d}a_{i}-\prod_{i=1}^{d}a^{\prime}_{i}|\leq d\cdot R^{d-1}\cdot\delta$
where this comes from mathematical induction. ∎
Appendix C Error Analysis of Visual Auto-Regressive Transformer
This section focuses on the error analysis of the $\operatorname{\mathsf{VAR}}$ model. In Section C.1, we introduce the Lipschitz property of a polynomial function. In Section C.2, we analyze the error propagation of inner product operation by giving two approximated inputs. In Section C.3, we analyze the error propagation of the fast attention $\mathsf{AAttC}(X)$. In Section C.4, we analyze the error between $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$ where $X^{\prime}$ is the approximated value of $X$. In Section C.5, we conduct an error analysis of the up-interpolation layer. In Section C.6, we conduct the error analysis of the $\operatorname{\mathsf{VAR}}$ transformer. Finally, in Section C.7, we conduct the error analysis of the Fast $\operatorname{\mathsf{VAR}}$ transformer.
C.1 Lipschitz of Polynomial
In this section, we introduce the Lipschitz property of polynomials.
Lemma C.1 (Lipschitz of polynomial).
Assuming the conditions below are satisfied:
  • Let $x\in\mathbb{R}$.
  • Let $x^{\prime}\in\mathbb{R}$ denote the approximated version of $x$.
  • Let $M>1$.
  • Suppose we have $|x|\leq M,|x^{\prime}|\leq M$.
  • Let $f$ be a polynomial with degree $g$.
We can then demonstrate that:
$|f(x)-f(x^{\prime})|\leq O(M^{g-1})\cdot|x-x^{\prime}|$ (2)
Proof.
Firstly, we can show
$f(x)=a_{g}x^{g}+\cdots+a_{1}x+a_{0}$
where for each $i\in[g]$, $a_{i}\in\mathbb{R}$.
Thus, we can show
$|f(x)-f(x^{\prime})|=$ $~{}|\sum_{i=1}^{g}a_{i}(x^{i}-x^{\prime i})|$
$\leq$ $~{}\sum_{i=1}^{g}|a_{i}(x^{i}-x^{\prime i})|$
$=$ $~{}\sum_{i=1}^{g}|a_{i}\cdot(x-x^{\prime})\cdot\sum_{j=0}^{i-1}x^{j}x^{\prime i-1-j}|$
The first step is derived from Eq. (2), the second from the triangle inequality, and the final step from simple algebra.
Then, we can move forward to show that
$~{}\sum_{i=1}^{g}|a_{i}\cdot(x-x^{\prime})\cdot\sum_{j=0}^{i-1}x^{j}x^{\prime i-1-j}|$
$\leq$ $~{}\sum_{i=1}^{g}|a_{i}\cdot(x-x^{\prime})|\cdot i\cdot M^{i-1}$
$=$ $~{}|x-x^{\prime}|\cdot\sum_{i=1}^{g}|a_{i}|\cdot i\cdot M^{i-1}$
$\leq$ $~{}O(M^{g-1})\cdot|x-x^{\prime}|$
The first step above is a consequence of the condition that $|x|\leq M$ and $|x^{\prime}|\leq M$, And the second derives from basic algebra, and the final step is due to $M>1$. ∎
C.2 Error Propagation of Inner Product
In this section, we conduct the error analysis of the inner product operation given two approximated inputs $u^{\prime}$ and $v^{\prime}$.
Lemma C.2 (Error propagation of inner product).
Assuming the conditions below are satisfied:
  • Let $u,v\in\mathbb{R}^{k}$ denote two vectors.
  • Define $u^{\prime},v^{\prime}\in\mathbb{R}^{k}$ as the approximation value of $u,v$.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let $\epsilon\in(0,0.1)$ denote the initial approximation error.
  • Suppose we have $\max\{\|u^{\prime}-u\|_{\infty},\|v^{\prime}-v\|_{\infty}\}\leq\epsilon$.
Then, we can prove that
$|\langle u^{\prime},v^{\prime}\rangle-\langle u,v\rangle|\leq 2k\epsilon M$
Proof.
Firstly, we can show that
$|\langle u^{\prime},v^{\prime}\rangle-\langle u,v\rangle|=$ $~{}|\sum_{i=1}^{k}(u^{\prime}_{1}v^{\prime}_{1}-u_{1}v_{1})|$
$\leq$ $~{}\sum_{i=1}^{k}|u^{\prime}_{1}v^{\prime}_{1}-u_{1}v_{1}|$
$\leq$ $~{}\sum_{i=1}^{k}|u^{\prime}_{1}(v^{\prime}_{1}-v_{1})+v_{1}(u^{\prime}_{1}-u_{1})|$
$\leq$ $~{}\sum_{i=1}^{k}(|u^{\prime}_{1}|\cdot|v^{\prime}_{1}-v_{1}|+|v_{1}|\cdot|u^{\prime}_{1}-u_{1}|)$
The first step results from simple algebra, the second from triangle inequality, the third from algebraic manipulation, and the final step again from triangle inequality.
Then, we can move forward to show
$~{}\sum_{i=1}^{k}(|u^{\prime}_{1}|\cdot|v^{\prime}_{1}-v_{1}|+|v_{1}|\cdot|u^{\prime}_{1}-u_{1}|)$
$\leq$ $~{}\sum_{i=1}^{k}2\cdot M\cdot\epsilon$
$\leq$ $~{}2k\epsilon M$
The first step is derived from the conditions that each entry is at most $M$, $\|u^{\prime}-u\|_{\infty}\leq\epsilon$ and $\|v^{\prime}-v\|_{\infty}\leq\epsilon$. The second step follows directly from algebraic manipulation.
C.3 Error Analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$
This section presents the error analysis between $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$.
Lemma C.3 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$, formal version of Lemma 5.4).
Assuming the conditions below are satisfied:
  • Let $X\in\mathbb{R}^{n\times d}$ denote the input matrix.
  • Define $X^{\prime}\in\mathbb{R}^{n\times d}$ as the approximation version of input matrix.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)$ as given in Definition 4.2.
  • Let $U,V\in\mathbb{R}^{n\times k}$ represent low-rank matrices constructed for polynomial approximation of attention matrix $\mathsf{AAttC}(X)$.
  • Let $f$ be a polynomial with degree $g$.
Then, we can show that
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X)\|_{\infty}\leq O(kM^{g+2}d)\cdot\epsilon$
Proof.
Firstly, given $X$ and $X^{\prime}$, we need to compute $Q,Q^{\prime},K,K^{\prime}$. And we demonstrate that
$\|Q-Q^{\prime}\|_{\infty}=$ $~{}\|X\cdot W_{Q}-X^{\prime}\cdot W_{Q}\|_{\infty}$
$=$ $~{}\|(\underbrace{X-X^{\prime}}_{n\times d})\cdot\underbrace{W_{Q}}_{d\times d}\|_{\infty}$
$\leq$ $~{}d\cdot\|X-X^{\prime}\|_{\infty}\cdot\|W_{Q}\|_{\infty}$
$\leq$ $~{}M\cdot d\cdot\epsilon$ (3)
The initial step is derived from the computation of matrix $Q$. The second step is a consequence of basic algebra, the third step arises from standard matrix multiplication, and the final step is a result of the condition $|X-X^{\prime}|_{\infty}\leq\epsilon$ and the fact that each entry is bounded by $M$.
In the same way, we can have $\|K-K^{\prime}\|_{\infty}\leq d\cdot\epsilon\cdot M$.
Then, we move forward to calculate $U\in\mathbb{R}^{n\times k}$ and $V\in\mathbb{R}^{n\times k}$. Specifically, for every $i\in[n]~{}\mathrm{and}~{}j\in[k]$, we have $U_{i,j}=f(Q_{i,1},\dots,Q_{i,d})~{}\mathrm{and}~{}V_{i,j}=f(K_{i,1},\dots,K_{i,d})$. Then, we can show that
$\|U-U^{\prime}\|_{\infty}\leq$ $~{}O(M^{g-1})\cdot\|Q-Q^{\prime}\|_{\infty}$
$\leq$ $~{}O(M^{g}d)\cdot\epsilon$
The first step above is derived from Lemma C.1 and the condition that each entry is bounded by $M$, while the second step results from Eq. (C.3) and basic algebra.
In the same way, we can have $\|V-V^{\prime}\|_{\infty}\leq O(M^{g}d)\cdot\epsilon$.
Finally, we can move forward to calculate $\mathsf{AAttC}(X^{\prime})=U^{\prime}V^{\prime\top}$ and $\mathsf{AAttC}(X)=UV^{\top}$. Then, for each $i\in[n],j\in[n]$, it can be demonstrated that
$|\mathsf{AAttC}(X^{\prime})_{i,j}-\mathsf{AAttC}(X)_{i,j}|=$ $~{}|\langle U^{\prime}_{i,*},V^{\prime}_{*,j}\rangle-\langle U_{i,*},V_{*,j}\rangle|$
$\leq$ $~{}2k\cdot O(M^{g}d)\cdot\epsilon\cdot M$
$\leq$ $~{}O(kM^{g+1}d)\cdot\epsilon$
The first step above is a result of basic algebra, the second step comes from Lemma C.2 and the lemma’s condition, and the final step is derived from basic algebra.
Thus, using the definition of the $\ell_{\infty}$ norm of a matrix, we can demonstrate that
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
Thus, we complete the proof. ∎
C.4 Error Analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$
In this section, we conduct the error analysis between $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$.
Lemma C.4 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$, formal version of Lemma 5.5).
Assuming the conditions below are satisfied:
  • Let $X\in\mathbb{R}^{n\times d}$ denote the input matrix.
  • Define $X^{\prime}\in\mathbb{R}^{n\times d}$ as the approximation version of input matrix.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the attention matrix $\mathsf{Attn}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 3.3.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 4.2.
  • Let $U,V\in\mathbb{R}^{n\times k}$ be low-rank matrices constructed for polynomial approximation of attention matrix $\mathsf{AAttC}(X)$.
  • Let $f$ be a polynomial with degree $g$.
We can demonstrate the following:
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{Attn}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
Proof.
It can be shown that
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{Attn}(X)\|_{\infty}=$ $~{}\|(\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X))+(\mathsf{AAttC}(X)-\mathsf{Attn}(X))\|_{\infty}$
$\leq$ $~{}\|(\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X))\|_{\infty}+\|(\mathsf{AAttC}(X)-\mathsf{Attn}(X))\|_{\infty}$
$\leq$ $~{}O(kM^{g+1}d)\cdot\epsilon+\epsilon$
$=$ $~{}O(kM^{g+1}d)\cdot\epsilon$
The first step is based on simple algebra, the second step is derived using the triangle inequality, the third step is obtained from Lemma C.3 and Lemma E.2, and the final step results from basic algebra.
C.5 Error Analysis of Up-Interpolation Layer
Furthermore, we still need the error analysis of up-interpolation layers.
Lemma C.5 (Error Analysis of Up-Interpolation Layer).
Assuming the conditions below are satisfied:
  • Let $X\in\mathbb{R}^{h\times w\times d}$ denote the input feature map.
  • Define $X^{\prime}\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times d}$ as the approximated input feature map.
  • Let $\phi_{\rm up}:\mathbb{R}^{h\times w\times d}\to\mathbb{R}^{h^{\prime}\times w^{\prime}\times d}$ represent the up-interpolation layer defined in Definition 3.2.
  • Let $W:\mathbb{R}\to\mathbb{R}$ be a bicubic spline kernel as defined in 3.1.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Let $\|X-X^{\prime}\|_{\infty}\leq\epsilon$.
Then we have
$\|\phi_{\rm up}(X)-\phi_{\rm up}(X^{\prime})\|_{\infty}\leq 16\epsilon$
Proof.
For each $i\in[h^{\prime}],j\in[w^{\prime}],l\in[d]$, we have
$|\phi_{\rm up}(X)_{i,j,l}-\phi_{\rm up}(X^{\prime})_{i,j,l}|=$ $~{}|\sum_{s=-1}^{2}\sum_{t=-1}^{2}W(s)\cdot(X_{\frac{ih}{h^{\prime}}+s,\frac{jw}{w^{\prime}}+t,l}-X^{\prime}_{\frac{ih}{h^{\prime}}+s,\frac{jw}{w^{\prime}}+t,l})\cdot W(t)|$
$\leq$ $~{}\sum_{s=-1}^{2}\sum_{t=-1}^{2}|W(s)\cdot(X_{\frac{ih}{h^{\prime}}+s,\frac{jw}{w^{\prime}}+t,l}-X^{\prime}_{\frac{ih}{h^{\prime}}+s,\frac{jw}{w^{\prime}}+t,l})\cdot W(t)|$
$\leq$ $~{}\sum_{s=-1}^{2}\sum_{t=-1}^{2}|W(s)\cdot W(t)|\cdot\epsilon$
$\leq$ $~{}16\epsilon$
The first step is based on Definition 3.2, the second step is derived using the triangle inequality, the third step is a consequence of $|X-X^{\prime}|_{\infty}\leq\epsilon$, and the final step follows from $W(x)\in[0,1]$ and basic algebra.
Then, according to the definition of the $l_{\infty}$ norm, we obtain
$\|\phi_{\rm up}(X)-\phi_{\rm up}(X^{\prime})\|_{\infty}\leq 16\epsilon$
Recall we have defined $\phi_{\mathrm{up}}:\mathbb{R}^{h\times w\times c}\to\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$ in Definition 3.2. Since there is no non-linear operation in $\phi_{\mathrm{up}}$, $\phi_{\mathrm{up}}$ is equivalent to a matrix multiplication operation, where the dimension of the matrix is $\mathbb{R}^{h^{\prime}w^{\prime}\times hw}$. For simplicity, we view $\phi_{\mathrm{up}}$ as a $\mathbb{R}^{h^{\prime}w^{\prime}\times hw}$ dimension matrix in the following proofs.
Remark C.6 (Applying $\phi_{\mathrm{up}}$ on $X\in\mathbb{R}^{n\times d}$).
The actual input of $\operatorname{\mathsf{VAR}}$ Transformer Layer are $r$ input token maps, $X_{1}\in\mathbb{R}^{h_{1}\times w_{1}\times d},\ldots,X_{r}\in\mathbb{R}^{h_{r}\times w_{r}\times d}$. We denote them as $X\in\mathbb{R}^{n\times d}$, where $n:=\sum_{i=1}^{r}h_{i}w_{i}$. We denote $\phi_{\mathrm{up}}(X)\in\mathbb{R}^{n^{\prime}\times d}$ as applying $\phi_{\mathrm{up}}$ to each $X_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times d}$ for $i\in[r]$, where $n^{\prime}=\sum_{i=1}^{r}h_{i}^{\prime}w_{i}^{\prime}$.
C.6 Error Analysis for $\operatorname{\mathsf{VAR}}$ Transformer
Then, we move forward to show the error propagation analysis for one $\operatorname{\mathsf{VAR}}$ Transformer Layer.
Lemma C.7 (Error propagation analysis for one $\operatorname{\mathsf{VAR}}$ Transformer Layer, formal version of Lemma 5.6).
Assuming the conditions below are satisfied:
  • Let $X\in\mathbb{R}^{n\times d}$ denote the input data matrix.
  • Define $X^{\prime}\in\mathbb{R}^{n\times d}$ as the approximation version of $X$.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let the $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\mathrm{var}}$ as given in Definition 3.5.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\rm fvar}$ as given in Definition 3.6.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the attention matrix $\mathsf{Attn}(X)\in\mathbb{R}^{n\times n}$ be defined as Definition 3.3.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 4.2.
  • Let $U,V\in\mathbb{R}^{n\times k}$ be low-rank matrices constructed for polynomial approximation of attention matrix $\mathsf{AAttC}(X)$.
  • Let $f$ be a polynomial with degree $g$.
It can be shown that
$\|\mathsf{F}_{\rm fvar}(X^{\prime})-\mathsf{F}_{\mathrm{var}}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
Proof.
Firstly, we demonstrate that
$\|\mathsf{F}_{\rm fvar}(X^{\prime})-\mathsf{F}_{\mathrm{var}}(X)\|_{\infty}=$ $~{}\|\mathsf{AAttC}(\phi_{\rm up}(X^{\prime}))-\mathsf{Attn}(\phi_{\rm up}(X))\|_{\infty}$
$\leq$ $~{}O(kM^{g+1}d)\cdot 16\epsilon$
$\leq$ $~{}O(kM^{g+1}d)\cdot\epsilon$
The first step is derived from Definition 3.7 and Definition 3.6, the second step is obtained using Lemma C.5 and Lemma C.4, and the final step follows from basic algebra. ∎
C.7 Error Analysis for the Fast $\operatorname{\mathsf{VAR}}$ Transformer
we perform an error analysis of the Fast $\operatorname{\mathsf{VAR}}$ Transformer in this section.
Lemma C.8 (Error analysis of the Fast $\operatorname{\mathsf{VAR}}$ Transformer, formal version of Lemma 5.7).
Assuming the conditions below are satisfied:
  • Let $X_{0}\in\mathbb{R}^{1\times d}$ denote the initial input data matrix.
  • Let the $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\mathrm{var}}$ be defined as Definition 3.5.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\mathrm{fvar}}$ as given in Definition 3.6.
  • Let $r\in\mathcal{N}$ represent the total iteration of the Fast $\operatorname{\mathsf{VAR}}$ model.
  • For $i\in[r]$, let $T_{i}(X)$ denote the intermediate variable of the $\operatorname{\mathsf{VAR}}$ model, as defined in Definition 3.7.
  • For $i\in[r]$, let $\widetilde{T}_{i}(X)$ denote the intermediate variable of the fast $\operatorname{\mathsf{VAR}}$ model, as defined in Definition 3.8.
  • Define $T_{r}(X)\in\mathbb{R}^{O(n^{2})\times d}$ a the final output of the $\operatorname{\mathsf{VAR}}$ Transformer.
  • Define $\widetilde{T}_{r}(X)\in\mathbb{R}^{O(n^{2})\times d}$ as the final output of the Fast $\operatorname{\mathsf{VAR}}$ Transformer.
  • Assume each entry in the matrices can be represented using $O(\log n)$ bits.
  • Let $U,V\in\mathbb{R}^{n\times k}$ be low-rank matrices constructed for polynomial approximation of attention matrix $\mathsf{AAttC}(X)$.
  • Let $f$ be a polynomial with degree $g$.
Then, we can show that the error bound of the final output $\widetilde{T}_{r}(X)$ as
$\|\widetilde{T}_{r}(X_{0})-T_{r}(X_{0})\|_{\infty}\leq 1/\operatorname{poly}(n)$
Proof.
We can conduct math induction as the following:
Consider the first iteration. We can show that.
$\|\widetilde{T}_{1}(X_{0})-T_{1}(X_{0})\|_{\infty}=$ $~{}\|\mathsf{F}_{\mathrm{fvar}}(X_{0})-\mathsf{F}_{\mathrm{var}}(X_{0})\|_{\infty}$
$\leq$ $~{}\|\mathsf{AAttC}(\phi_{\rm up}(X_{0}))-\mathsf{Attn}(\phi_{\rm up}(X_{0}))\|_{\infty}$
$\leq$ $~{}1/\operatorname{poly}(n)$
The first step is based on Definition 3.8 and Definition 3.7, the second step is derived from Definition 3.6 and Definition 3.5, and the final step follows from Lemma E.2.
Assume that the following statement is true for the $k$-th iteration:
$\|\widetilde{T}_{k}(X_{0})-T_{k}(X_{0})\|_{\infty}\leq 1/\operatorname{poly}(n)$ (4)
Then we move forward to consider the $k+1$-th iteration as the following:
$\|\widetilde{T}_{k+1}(X_{0})-T_{k+1}(X_{0})\|_{\infty}=$ $~{}\|\mathsf{F}_{\mathrm{fvar}}(\widetilde{T}_{k}(X_{0}))-\mathsf{F}_{\mathrm{var}}(T_{k}(X_{0}))\|_{\infty}$
$\leq$ $~{}O(k\cdot(\operatorname{poly}(n))^{g+1}\cdot d)\cdot(1/\operatorname{poly}(n))$
$\leq$ $~{}1/\operatorname{poly}(n)$
The first step is based on Definition 3.8 and Definition 3.7, the second step is derived from Lemma C.7, the fact that each entry in the matrices can be represented using $O(\log(n))$ bits, and Eq. (4), while the final step results from basic algebra.
Finally, we can use math induction to show that
$\|\widetilde{T}_{r}(X_{0})-T_{r}(X_{0})\|_{\infty}\leq 1/\operatorname{poly}(n).$
Thus, we complete the proof. ∎
Appendix D Error Analysis of VQVAE Decoder
In this section, we conduct the error analysis of the VQ-VAE Decoder. Firstly, the following lemma presents the error analysis of the Convolution Layer.
Lemma D.1 (Error analysis of Convolution Layer).
Assuming the conditions below are satisfied:
  • Let $X\in\mathbb{R}^{h\times w\times c_{\rm in}}$ denote the input feature map.
  • Let $X^{\prime}\in\mathbb{R}^{h\times w\times c_{\rm out}}$ denote the output feature map.
  • Let $\phi_{\rm conv}:\mathbb{R}^{h\times w\times c_{\rm in}}\to\mathbb{R}^{h\times w\times c_{\rm out}}$ denote the convolution layer defined in Definition 3.9.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Let $\|X-X^{\prime}\|_{\infty}\leq\epsilon$.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let $C=9c_{\rm in}$ denote a constant.
Then we have
$\|\phi_{\rm conv}(X)-\phi_{\rm conv}(X^{\prime})\|_{\infty}\leq C\epsilon M$
Proof.
For each $i\in[h],j\in[w],l\in[c_{\rm out}]$, we have
$|\phi_{\rm conv}(X)_{i,j,l}-\phi_{\rm conv}(X^{\prime})_{i,j,l}|=$ $~{}|\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{c=1}^{c_{\rm in}}(X_{i+m-1,j+n-1,c}-X^{\prime}_{i+m-1,j+n-1,c})\cdot K^{l}_{m,n,c}|$
$\leq$ $~{}\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{c=1}^{c_{\rm in}}|(X_{i+m-1,j+n-1,c}-X^{\prime}_{i+m-1,j+n-1,c})\cdot K^{l}_{m,n,c}|$
$\leq$ $~{}\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{c=1}^{c_{\rm in}}\epsilon\cdot M$
$\leq$ $~{}9\cdot c_{\rm in}\epsilon M$
$=$ $~{}C\epsilon M$
The 1st step is a consequence of Definition 3.9, the 2nd step is based on the triangle inequality, the 3rd is a result of the lemma’s conditions, the fourth arises from elementary algebra, and the final step stems from the definition of $C$. ∎
Then, we can show the lemma, which presents the error analysis of the Fast VQ-VAE Decoder.
Lemma D.2 (Error analysis of Fast VQ-VAE Decoder ).
In the case that the following conditions are satisfied:
  • Let $X\in\mathbb{R}^{n\times d}$ denote the input matrix.
  • Let the up-interpolation Layer $\phi_{\rm up}$ be defined as Definition 3.2.
  • Let the convolution layer $\phi_{\rm conv}$ be defined as Definition 3.9.
  • Let the attention layer $\mathsf{Attn}$ be defined as Definition 3.4
  • Let the fast attention layer $\mathsf{AAttC}$ be defined as Definition 4.2.
  • Let the VQ-VAE Decoder be the composition of a constant number of up-interpolation layers, convolutions layers, and attention layers.
  • Let the Fast VQ-VAE Decoder be defined as substituting all $\mathsf{Attn}$ layers in VQ-VAE with $\mathsf{AAttC}$ layers.
Then, we can show that the approximation error of the Fast VQ-VAE Decoder can be bounded by $1/\operatorname{poly}(n)$.
Proof.
By Lemma C.7, we have shown that fast attention computation will introduce an approximation error no more than $1/\operatorname{poly}(n)$.
Then, by Lemma C.5 and Lemma D.1, we still can bounded have shown the approximation error by $1/\operatorname{poly}(n)$ after passing another up-interpolation layer, convolutions layer.
Since VQ-VAE is a composition of a constant number of up-interpolation layers, convolution layers, and attention layers, the overall approximation error can still be bounded by $1/\operatorname{poly}(n)$. ∎
Appendix E Running Time
In this section, we conduct the running time analysis of every component of the $\operatorname{\mathsf{VAR}}$ model and the fast $\operatorname{\mathsf{VAR}}$ model. In Section E.1, we conduct the running time analysis of the $\operatorname{\mathsf{VAR}}$ transformer and the fast transformer. In Section E.2, we conduct the running time analysis of the feature map reconstruction block. In Section E.3, we conduct the running time analysis of the VQVAE Decoder and fast VQVAE Decoder.
E.1 Phase 1: Running Time of Token Maps Generation
In this section, we present a lemma on the time complexity of the token maps generation phase of $\operatorname{\mathsf{VAR}}$ and reference the results in [5] to derive a faster variant of $\operatorname{\mathsf{VAR}}$.
Lemma E.1 (Running time of $\operatorname{\mathsf{VAR}}$ Transformer).
Assuming the conditions below are satisfied:
  • Let $K$ denote the total number of the token maps.
  • Let $k\in[K]$.
  • Let $d$ denote the embedding size of each token.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k\in[K]$, the $k$-th token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$.
  • Let $r_{K}\in\mathbb{R}^{n\times n\times d}$ denote the last scale token map, where $n=\alpha^{K-1}$.
then the time complexity of $\operatorname{\mathsf{VAR}}$ generation is $O(n^{4}d)$.
Proof.
Firstly, we consider the $k$-th autoregressive token map generation. We use $L_{k}$ to denote the total number of the tokens input into the $\operatorname{\mathsf{VAR}}$ transformer at $k$-th step and $L_{k}$ can be calculated as the following:
$L_{k}=$ $~{}\sum_{i=1}^{k}(\alpha^{i-1})^{2}$
$=$ $~{}\frac{\alpha^{2k}-1}{\alpha^{2}-1}$
$\leq$ $~{}\frac{\alpha^{2k}}{\alpha^{2}-1}$
$\leq$ $~{}\frac{\alpha^{2k}}{0.5\alpha^{2}}$
$=$ $~{}2\cdot\alpha^{2k-2}$ (5)
In the first step, we use the condition of this lemma. The second and third steps are a consequence of basic algebra. The fourth step is due to $\alpha\geq 2$, and the last step is derived from elementary algebra.
Thus, at the $k$-th step, we use $X_{k}\in\mathbb{R}^{L_{k}\times d}$ to denote the input matrix. The transformer computation cost at the $k$-th step is $O(L_{k}^{2}d)$. We then sum up the computation time across all $K$ steps:
${\cal T}=$ $~{}O((L_{1}^{2}+L_{2}^{2}+\dots+L_{K}^{2})\cdot d)$
$\leq$ $~{}O(\sum_{k=1}^{\log_{a}{n}+1}(2\cdot\alpha^{2k-2})^{2}\cdot d)$
$=$ $~{}O(\sum_{k=1}^{\log_{a}{n}+1}4\alpha^{4k-4}\cdot d)$
$=$ $~{}O(n^{4}d)$
In the first step, the total time is obtained by summing the times for each $k$-th step, while the second step is a consequence of Eq. (E.1), and the final step is attributed to elementary algebra.
Then, we show a lemma that demonstrates a fast way to compute attention in [5].
Lemma E.2 (Fast Attention Computation, Theorem 1.4 of [5]).
Let $\mathsf{AAttC}$ be defined as Definition 4.2. Then we have $\mathsf{AAttC}(n,d=\log n,R=(\sqrt{\log n}),\delta=1/\operatorname{poly}(n))$ can be solved in time $\mathcal{T}_{\rm mat}(n,n^{o(1)},d)=n^{1+o(1)}$.
Now we can apply the result Lemma E.2 to the $\operatorname{\mathsf{VAR}}$ Transformer.
Lemma E.3 (Running time of Fast $\operatorname{\mathsf{VAR}}$ Transformer, formal version of Lemma 5.1).
Assuming the conditions below are satisfied:
  • Denote $K$ as the total number of generated token maps.
  • Let $k\in[K]$.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k\in[K]$, the $k$-th token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$.
  • Let $r_{K}\in\mathbb{R}^{n\times n\times d}$ denote the last scale token map, where $n=\alpha^{K-1}$.
  • Let $d=O(\log(n))$ denote embedding size of each token.
Then, the total running time of the $\operatorname{\mathsf{VAR}}$ Transformer for generating token maps can be accelerated to $O(n^{2+o(1)}d)$.
Proof.
To generate the $k$-th token map, let the input of this step be $X\in\mathbb{R}^{L_{k}\times d}$. And we have
$L_{k}\leq 2\cdot\alpha^{2k-2}$ (6)
where this step is a consequence of Eq. (E.1). So the transformer computation cost at the $k$-th step can be improved from $O(L_{k}^{2}d)$ to $O(L_{k}^{1}d)$ by using the result of Lemma E.2. Then, we sum up the computation time across all $K$ steps:
${\cal T}=$ $~{}O((L_{1}+L_{2}+\dots+L_{K})\cdot d)$
$\leq$ $~{}O(\sum_{k=1}^{\log_{\alpha}{n}+1}(2\cdot\alpha^{2k-2})^{1+o(1)}\cdot d)$
$=$ $~{}O(n^{2+o(1)}d)$
The computation time in the first step is determined by summing all $k$-th step times, while the second step is a consequence of Eq. (6), and the last step is attributed to elementary algebra. ∎
E.2 Phase 2: Running Time of Feature Map Reconstruction
In this section, we analyze the running time of the feature map reconstruction layer.
Lemma E.4 (Running time of Feature Map Reconstruction Layer, formal version of Lemma 5.2).
Assuming the conditions below are satisfied:
  • Define $K$ as the total number of the token maps.
  • Let $k\in[K]$.
  • Let $d$ denote the embedding size of each token.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k\in[K]$, the $k$-th token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$.
  • Let $r_{K}\in\mathbb{R}^{n\times n\times d}$ denote the last scale token map, where $n=\alpha^{K-1}$.
then the total runtime of the $\operatorname{\mathsf{VAR}}$ models for feature map reconstruction is $O(n^{2}d^{2}\log n)$.
Proof.
For each $k\in[K]$, $\operatorname{\mathsf{VAR}}$ Model will up-interpolate token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$ to $r^{\prime}_{k}\in\mathbb{R}^{n\times n\times d}$ by using bicubic interpolation defined in Definition 3.2. Specifically, the computation of each entry of $r^{\prime}_{k}$ requires $32$ multiplications and $16$ additions (see more details in Definition 3.2). Thus, the computation cost for the up-interpolation of each token map is
${\cal T}^{k}_{\rm up}=$ $~{}48\cdot n^{2}d$
$=$ $~{}O(n^{2}d)$
There are total $O(\log n)$ token maps needed to be up-interpolated, so the total time for up-interpolation is
${\cal T}_{\rm up}=$ $~{}O(n^{2}d)\cdot O(\log n)$
$=$ $~{}O(n^{2}d\log n)$
The computation time in the first step is determined by summing the $\log(n)$ up-interpolation time, and the second step is a consequence of elementary algebra.
Furthermore, to address the information loss in the up-interpolation process, the $\operatorname{\mathsf{VAR}}$ Model uses a convolution operation $\phi(\cdot)$ on the token map $\{r^{\prime}_{1},\dots,r^{\prime}_{K}\}$ generated by up-interpolation. We assume the convolution kernel size is $3\times 3\times d$, and the convolution layer does not change the dimension of each token map, i.e., for each $i\in[K]$, $\phi(r^{\prime}_{i})\in\mathbb{R}^{n\times n\times d}$. Hence, for every entry in $\phi(r^{\prime}_{i})$, it needs $O(d)$ operations. Then, we can have the convolution computation time for one token map is
${\cal T}^{k}_{\rm conv}=$ $~{}O(d)\cdot n^{2}d$
$=$ $~{}O(n^{2}d^{2})$
In the first step, the total computation time is obtained by adding the times for the $n\times n\times d$ entries, while the second step results from simple algebra.
There are total $O(\log n)$ token maps needed to be passed to the convolution layer, so the total time for convolution operations is
${\cal T}_{\rm conv}=$ $~{}O(\log n)\cdot O(n^{2}d^{2})$
$=$ $~{}O(n^{2}d^{2}\log n)$
In the first step, the total time is obtained by $O(\log(n))$ convolution operations, while the second step results from simple algebra.
Then, the $\operatorname{\mathsf{VAR}}$ Model will sum up $O(\log n)$ token maps processed by convolution layers, where each token map has a size of $n\times n\times d$. Thus, the computation cost of addition needs
${\cal T}_{\rm add}=$ $~{}O(\log n)\cdot(n^{2}d)$
$=$ $~{}O(n^{2}d\log n)$
In the first step, token maps are added element-wise, and there are $O(\log(n))$ token maps in total, while the second step results from basic algebra.
Hence, the running time of feature map reconstruction is as follows:
${\cal T}_{\rm rc}=$ $~{}{\cal T}_{\rm up}+{\cal T}_{\rm conv}+{\cal T}_{\rm add}$
$=$ $~{}O(n^{2}d^{2}\log n)$
The first step is derived by summing the times for up-interpolation operations, convolution operations, and token map additions, while the second step is due to basic algebra. ∎
E.3 Phase 3: Running Time of VQ-VAE Decoder
In this section, we analyze the running time of the VQ-VAE Decoder and fast VQ-VAE Decoder.
Lemma E.5 (Running time of VQ-VAE Decoder).
Assuming the conditions below are satisfied:
  • Let $k_{1},k_{2},k_{3}\in\mathbb{N}$ be constant numbers.
  • Given $X\in\mathbb{R}^{n\times n\times d}$ as the input feature map.
  • Assume that there are $k_{1}$ up-interpolation layers $\phi_{\rm up}$ defined in Definition 3.2.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{O(h)\times O(w)\times d}$.
  • We assume there are $k_{2}$ attention layers $\mathsf{Attn}$ defined in Definition 3.4.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, the $i$-th attention layer’s output $\mathsf{Attn}(M)\in\mathbb{R}^{h\times w\times d}$.
  • We assume there are $k_{3}$ convolution layers $\phi_{\rm conv}$ defined in Definition 3.9.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
then the total running time of the VQ-VAE Decoder is $O(n^{4}d)$.
Proof.
By the condition, we can have that for each $l\in[k_{1}+k_{2}+k_{3}]$, the size of the output $M^{l}$ of any intermediate layer (up-interpolation layer, convolution layer, attention layer) is $O(n)\times O(n)\times O(d)$. The running time can be computed as follows:
Part 1. Running time of Up-interpolation layers. For each $l\in[k_{1}]$, we assume $M^{l}\in\mathbb{R}^{O(n)\times O(n)\times O(d)}$ as the output feature map from the $l$-th up-interpolation layer. For every entry of $M^{l}$, it needs $32$ multiplications and $16$ additions (see more details in Definition 3.2). Thus, the computation cost for the feature map $M^{l}$ is
${\cal T}^{l}_{\rm up}=$ $~{}48\cdot O(n)\cdot O(n)\cdot O(d)$
$=$ $~{}O(n^{2}d)$
The first step is derived by summing the computation costs for each entry of $M^{l}$, while the second step is due to basic algebra.
Since there are $k_{1}$ up-interpolation layers in total, the total time of the up-interpolation layers in the VQ-VAE decoder is
${\cal T}_{\rm up}=$ $~{}k_{1}\cdot O(n^{2}d)$
$=$ $~{}O(n^{2}d)$ (7)
The first step is derived by summing the computation costs for each up-interpolation layer, while the second step is due to the fact that $k_{1}$ is a constant number.
Part 2. Running time of Attention layers. For each $l\in[k_{2}]$, we assume $M^{l-1}\in\mathbb{R}^{O(n)\times O(n)\times O(d)}$ as the feature map used as input for the $l$-th attention layer. We can consider the input of this attention layer as a sequence with length $O(n^{2})$ and embedding dimension $O(d)$. Hence, the computation cost of this attention layer is
${\cal T}^{l}_{\rm attn}=O(n^{4}d)$
Since there are $k_{2}$ attention layers in total, the total time of the attention layers in the VQ-VAE decoder is
${\cal T}_{\rm attn}=$ $~{}k_{2}\cdot O(n^{4}d)$
$=$ $~{}O(n^{4}d)$
The first step is derived by summing the computation costs for each up-interpolation layer, while the second step is due to the fact that $k_{2}$ is a constant.
Part 3. Running time of Convolution layers. For each $l\in[k_{3}]$, we assume $M^{l}\in\mathbb{R}^{O(n)\times O(n)\times O(d)}$ as the output feature map of the $l$-th convolution layer. For every entry of $M^{l}$, it needs $O(d)$ operations. Thus, the computation cost for the feature map $M^{l}$ is
${\cal T}^{l}_{\rm conv}=$ $~{}O(d)\cdot O(n^{2}d)$
$=$ $~{}O(n^{2}d^{2})$
The first step is derived by summing the computation costs for each entry of $M^{l}$, while the second step stems from basic algebra.
Since there are $k_{3}$ convolution layers in total, the total time of the convolution layers in the VQ-VAE decoder is
${\cal T}_{\rm conv}=$ $~{}k_{3}\cdot O(n^{2}d^{2})$
$=$ $~{}O(n^{2}d^{2})$ (8)
The first step is derived by summing the computation costs for each convolution layer, while the second step is due to the fact that $k_{3}$ is a constant.
Finally, the computation cost of the VQ-VAE decoder can be calculated as follows:
${\cal T}_{\rm dec}=$ $~{}{\cal T}_{\rm up}+{\cal T}_{\rm attn}+{\cal T}_{\rm conv}$
$=$ $~{}O(n^{4}d)$
The first step results from summing the computation costs of up-interpolation, attention, and convolution layers, while the second step is a consequence of $n\gg d$.
Then, we move forward to show the running time of the fast VQ-VAE decoder.
Lemma E.6 (Running time of Fast VQ-VAE Decoder, formal version of Lemma 5.3).
In the case that the following conditions are satisfied:
  • Assume $k_{1},k_{2},k_{3}\in\mathbb{N}$ are constant numbers.
  • We represent the input feature map by $X\in\mathbb{R}^{n\times n\times d}$.
  • We assume there are $k_{1}$ up-interpolation layers $\phi_{\rm up}$ defined in Definition 3.2.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{O(h)\times O(w)\times d}$.
  • We assume there are $k_{2}$ attention layers $\mathsf{Attn}$ defined in Definition 3.4.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, the $i$-th attention layer’s output $\mathsf{Attn}(M)\in\mathbb{R}^{h\times w\times d}$.
  • We assume there are $k_{3}$ convolution layers $\phi_{\rm conv}$ defined in Definition 3.9.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
then the total runtime of the VQ-VAE decoder can be accelerated to $O(n^{2+o(1)}d)$.
Proof.
As the same in Eq. (E.5) and Eq. (E.5), the computation cost for up-interpolation layers and convolution layers in VQ-VAE decoder still needs $O(n^{2}d)$ and $O(n^{2}d^{2})$, respectively.
For each $l\in[k_{2}]$, we assume $M^{l-1}\in\mathbb{R}^{O(n)\times O(n)\times O(d)}$ as the input feature map for the $l$-th attention layer. We can consider the input of the attention layer as a sequence with length $O(n^{2})$ and embedding dimension $O(d)$. By using the result of Lemma E.2, the computation cost of $M^{l}$ can be speed up to
${\cal T}^{l}_{\rm attn}=O(n^{2+o(1)}d)$
Since there are $k_{2}$ attention layers in total, the total computation cost of the attention layers in the VQ-VAE decoder is
${\cal T}_{\rm attn}=$ $~{}k_{2}\cdot O(n^{2+o(1)}d)$
$=$ $~{}O(n^{2+o(1)}d)$
The computation cost in the first step is obtained by adding the costs of the up-interpolation layers, attention layers, and convolution layers, while the second step stems from $k_{2}$ is a constant.
Thus, the total runtime of the VQ-VAE decoder can be calculated as follows:
${\cal T}_{\rm dec}=$ $~{}{\cal T}_{\rm up}+{\cal T}_{\rm attn}+{\cal T}_{\rm conv}$
$=$ $~{}O(n^{2+o(1)}d)$
The computation cost in the first step is obtained by adding the costs of the up-interpolation layers, attention layers, and convolution layers, while the second step comes from $n\gg d$. ∎